Ivanhoe provides shuttle service between four hotels near a medical center and an international airport. Ivanhoe uses two 10 passengers vans to offer 12 round trips per day. A recent months activity in the form of a cost-volume-profit income statement is shown below.

Sales (1,300 passengers).   $26,000
Variable Costs
Fuel $6,334
Tolls and Parking $1,685
Maintenance $1,081.         $9,100
Contribution Margin.           $16,900

Fixed Costs
Salaries    $12,900
Depreciation $1,250
Insurance    $930
Net Income.     $1,820

Calculate the break-even point in dollars.

Calculate the break-even point in numbers of Passengers.

Without calculations, determine the contribution margin at the break-even point.
Answer #1

Contribution margin ratio = 16900/26000 = 65%

a) Break even sales = Fixed cost/Contribution margin ratio = 15080/.65 = $23200

b) Break even unit = 23200/20 = 1160 Passengers

c) Contribution margin at break even = Fixed cost = $23200